OpenStudy (anonymous):

evaluate the following expression for x=5 and y=-7 5x+7y-3x+8

OpenStudy (anonymous):

5(5)+7(-7)-3(5)+8=31

OpenStudy (anonymous):

oops typo. it comes out to -31 not 31

OpenStudy (anonymous):

5(5)+7(-7)-3(5)+8=25-49-15+8=-31